OpenStudy (anonymous):

i hv jst started with leibnitz`s theorem and cant understand it plz someone explane me this one

OpenStudy (anonymous):

if \[y= x ^{2}e ^{x}\] then show that \[y _{n} = y _{2}\frac{ n(n-1) }{ 2 } - n(n-2) y _{2} + \frac{ (n-1)(n-2) }{ 2 }y\]

OpenStudy (amistre64):

we might want to start with defining what you think the leibniz thrm says ...

OpenStudy (anonymous):

i know the formula and its method i hv done just a few questions which dealt with just expanding or solving them by applying the formula this is the first of show that kind... and i am not getting the answer

OpenStudy (amistre64):

this one? \[\sum_{k=0}^{n}~\binom{n}{k}\frac{d^k}{dx^k}[u(x)]~\frac{d^{n-k}}{dx^{n-k}}[v(x)]\]

OpenStudy (amistre64):

or do we have to brute it out to find a recursive pattern that would end up being that?

OpenStudy (amistre64):

\[x^2e^x\]\[2xe^x+x^2e^x\]\[2e^x+4xe^x+x^2e^x\]\[6e^x+6xe^x+x^2e^x\]\[12e^x+8xe^x+x^2e^x\] etc...

OpenStudy (anonymous):

yes that was of this kind yes

OpenStudy (amistre64):

so successive product rules ...

OpenStudy (amistre64):

\[f = 0e^x+0xe^x+x^2e^x\] \[f^1 = 0e^x+2xe^x+x^2e^x\] \[f^2 = 2e^x+4xe^x+x^2e^x\] \[f^3 = 6e^x+6xe^x+x^2e^x\] \[f^4 = 12e^x+8xe^x+x^2e^x\] \[f^5 = 20e^x+10xe^x+x^2e^x\] ... it appears to be this pattern to me \[f^n=n(n-1)e^x+2n~xe^x+x^2e^x\]

OpenStudy (amistre64):

can we conform that to the setup you have?

OpenStudy (amistre64):

looks like someone else came to the same conclusion ... Example 7.4.2 is similar except a=1 in your case http://wdjoyner.com/teach/calc1-sage/html/node106.html